# Morita equivalence of Nijenhuis structures

### v1

@@ 4.1 Quasi-symplectic-Nijenhuis groupoids
Let $G_{0}\leftarrow(P_{1},\varpi_{1},J_{1})\to H_{0}$ be a Morita equivalence between $(\mathscr{G},\omega_{1},\eta_{1},N_{1})$ and $(\mathscr{H},\omega_{2},\eta_{2},N_{2})$. Similarly, let $H_{0}\leftarrow(P_{2},\varpi_{2},J_{2})\to G_{0}^{\prime}$ be a Morita equivalence between $(\mathscr{H},\omega_{2},\eta_{2},N_{2})$ and $(\mathscr{G}^{\prime},\omega_{3},\eta_{3},N_{3})$. Let $P=P_{1}\diamond P_{2}$ be the quotient of the fiber product $P_{1}\times_{H_{0}}P_{2}$ by the diagonal action of $\mathscr{H}$, and denote by $\pi:P_{1}\times_{H_{0}}P_{2}\to P$ the natural projection. Our proof relies on the following results:
  1. 1.
    Theorem 2.11 states that $(P,J)$ defines a Morita bibundle between the Nijenhuis groupoids $(\mathscr{G},N_{1})$ and $(\mathscr{G}^{\prime},N_{3})$, where the Nijenhuis tensor field $J\in\Omega^{1}(P,TP)$ being the projection of $J_{1}\times J_{2}$.
  2. 2.
    From [41, Theorem 4.5], we know that $(P,\varpi)$ defines a Morita bibundle between the quasi-symplectic groupoids $(\mathscr{G},\omega_{1},\eta_{1})$ and $(\mathscr{G}^{\prime},\omega_{3},\eta_{3})$, where the $2$-form $\varpi\in\Omega^{2}(P)$ defined by equation (41)
It remains only to prove the compatibility between the $(1,1)$-tensor field $J$ and the $2$-form $\varpi$ on the quotient manifold $P$.
Compatibility on the fiber product. First, we show that the compatibility holds for the $(1,1)$-tensor field $J_{1}\times J_{2}\in\Omega^{1}\left(P_{1}\times_{H_{0}}P_{2},T(P_{1}\times P_{2})\right)$ and the $2$-form $\pi^{*}\varpi=\mathrm{pr}_{1}^{*}\varpi_{1}+\mathrm{pr}_{2}^{*}\varpi_{2}\in\Omega^{2}\left(P_{1}\times_{H_{0}}P_{2}\right)$. To verify the first condition, Equation (42), let $(v_{1},v_{2})$ be a vector field on $P_{1}\times_{H_{0}}P_{2}$. Using the fact that $(\varpi_{1},J_{1})$ and $(\varpi_{2},J_{2})$ are compatible pairs, we have
@@
$\mathrm{d}\left(\left(\pi^{*}\varpi\right)_{J_{1}\times J_{2}}\right)=\left(\mathrm{d}\left(\pi^{*}\varpi\right)\right)_{J_{1}\times J_{2}}.$ (44)
Thus, the pair $(\pi^{*}\varpi,\times J_{2})$ is compatible.
Descent to the quotient. Finally, these compatibility relations descend to the quotient $P$. Consider a vector field $\bar{v}\in\mathfrak{X}(P)$ as the image under $\mathrm{d}\pi$ of some $\pi$-projectable vector field $v=(v_{1},v_{2})\in\mathfrak{X}\left(P_{1}\times_{H_{0}}P_{2}\right)$. Recall that $J_{1}\times J_{2}$ preserves projectable vector fields and that it is $\pi$-related to $J$. Using (41) and evaluating at $\bar{v}=\mathrm{d}\pi(v)$, we find:
$\varpi^{\flat}\circ J\left(\mathrm{d}\pi(v)\right)$ $=\varpi^{\flat}\circ\mathrm{d}\pi\left(J_{1}\times J_{2}\left(v\right)\right)$
@@
$=(J_{1}\times J_{2})^{*}\left(\mathrm{pr}_{1}^{*}\varpi_{1}+\mathrm{pr}_{2}^{*}\varpi_{2}\right)^{\flat}(v)$
@@
$=J^{*}\circ\varpi^{\flat}(\mathrm{d}\pi(v)).$
Therefore $\varpi^{\flat}\circ J=J^{*}\circ\varpi^{\flat}$. Similarly, using (44) and the naturality of the exterior derivative, we obtain:
$\left(\mathrm{d}\varpi\right)_{J}\left(\mathrm{d}\pi(u),\mathrm{d}\pi(v),\mathrm{d}\pi(w)\right)$ $=\mathrm{d}\varpi\left(J\left(\mathrm{d}\pi(u)\right),\mathrm{d}\pi(v),\mathrm{d}\pi(w)\right)$
@@
$=\mathrm{d}\left(\varpi_{J}\right)\left(\mathrm{d}\pi(u),\mathrm{d}\pi(v),\mathrm{d}\pi(w)\right).$
Therefore $\mathrm{d}(\varpi_{J})=(\mathrm{d}\varpi)_{J}$. This completes the proof. ∎
4.2 Twisted Dirac-Nijenhuis structures
@@
$\llbracket(u,\alpha),(v,\beta)\rrbracket_{\eta}\coloneqq\left(\left[u,v\right],\mathscr{L}_{u}\beta-\iota_{v}\mathrm{d}\alpha+\iota_{u\wedge v}\eta\right),\quad\eta\in\Omega^{3}(M).$ (45)
Every twisted Dirac structure $L$ is intrinsically a Lie algebroid, where the anchor is the projection to $TM$ and the bracket is the restriction of the $\eta$-Courant bracket. The conditions for when an arbitrary Lie algebroid is isomorphic to a Dirac structure are detailed in [6].
Integrating Dirac structures.
The precise correspondence between twisted Dirac structures are quasi-symplectic groupoids is established in [6, Theorem 2.2] and [6, Theorem 2.4]: an integrable Lie algebroid $A_{\mathscr{G}}$ is isomorphic to an $\eta$-twisted Dirac structure $L$ on its base if and only if its source-simply connected integration $\mathscr{G}$ admits a quasi-symplectic structure. This correspondence requires the target map to be a forward Dirac map. A map $\mu:(M,L_{M})\to(N,L_{N})$ is said to be a forward Dirac map if
$(L_{N})_{\mu(p)}\coloneqq\left\{(\mathrm{d}\mu(v),\alpha)\in\mathbb{T}_{\mu(p)}N\mid v\in T_{p}M,(v,(\mathrm{d}\mu)^{*}(\alpha))\in L_{M}\right\}.$ (46)
Twisted Dirac-Nijenhuis structures.
@@
We recall from [8] a generalization of the compatibility for pairs given in Equations (42) and (43). Let $L\subseteq TM\oplus T^{*}M$ be a subbundle. A $(1,1)$-tensor field $r\in\Omega^{1}(M,TM)$ is said to be compatible with $L$ if
$\left(r,r^{*}\right)(L)\subseteq L\qquad\text{and}\qquad\mathbb{D}^{r}_{v}\left(\Gamma(L)\right)\subseteq\Gamma(L),$ (47)
@@
Example 4.8 (Poisson-Nijenhuis manifold [28, 23]).
@@
$\pi^{\sharp}\circ r^{*}=r\circ\pi^{\sharp}$ (48)
$R_{\pi}^{r}\left(v,\alpha\right)\coloneqq\pi^{\sharp}\left(\mathscr{L}_{v}r^{*}(\alpha)-\mathscr{L}_{r(v)}\alpha\right)-\left(\mathscr{L}_{\pi^{\sharp}(\alpha)}r\right)(v)=0,$ (49)
@@
The relation between $(1,1)$-tensor fields $r\in\Omega^{1}(M,TM)$ compatible with a Dirac structure $L\subseteq\mathbb{T}M$ and the Lie algebroid defined by $L$ is provided by [8, Lemma 6.1]: if $r\in\Omega^{1}(M,TM)$ is compatible with $L$, then the $1$-derivation $\left(\mathbb{D}^{r}|_{\Gamma(L)},(r,r^{*}),r\right)$ is a IM $1$-derivation on the Lie algebroid $L$. Moreover, if $r$ is Nijenhuis then the corresponding $1$-derivation is Nijenhuis.
@@
See [8, Section 3.3]. ∎
Integrating Dirac-Nijenhuis structures.
The following theorem is an adaptation of [8, Theorem 6.3] to the twisted case where $\eta\neq 0$. The proof is similar.
@@
Let $(\mathscr{G},\omega,\eta)$ be a quasi-symplectic groupoid integrating the $\eta$-twisted Dirac structure $L\subseteq\mathbb{T}M$. Let $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ be a Nijenhuis $(1,1)$-tensor field integrating the $1$-derivation $\left(D,\ell,r\right)$. If $\omega$ and $N$ are compatible (Equations (42) and (43)), then $L$ and $r$ are compatible in the sense of Equations (47), and the converse holds when $\mathscr{G}$ is source connected.
@@
4.3 Global and infinitesimal Morita equivalences of compatible structures
In this subsection, we introduce a notion of Morita equivalence for Dirac-Nijenhuis structures which corresponds to the infinitesimal counterpart of Morita equivalenence for quasi-symplectic Nijenhuis groupoids. To do so, we first need to review the Morita equivalence of Dirac manifolds themselves. Similarly to how a Morita equivalence of Poisson manifolds can be expressed in terms of symplectic realizations, Morita equivalences of Dirac structures are expressed in terms of the so-called Dirac realizations, or more precisely a dual pair, a concept detailed in [17].
In the following, we fix two twisted Dirac manifolds $(M_{1},L_{1},\eta_{1})$ and $(M_{2},L_{2},\eta_{2})$ and consider a manifold $P$ equipped with a $2$-form $\varpi\in\Omega^{2}(P)$. Consider surjective submersions $\mu_{1}:P\to M_{1}$ and $\mu_{2}:P\to M_{2}$, define $\eta\coloneqq\mu_{2}^{*}\eta_{2}-\mu_{1}^{*}\eta_{1}$, and equip $P$ with the $\eta$-twisted Dirac structure $L_{\varpi}$. We denote the opposite Dirac structure on $M_{2}$ by $\bar{L}_{2}\coloneqq\{(v,-\alpha)\mid(v,\alpha)\in L_{2}\}$. Furthermore, let $\mathscr{a}_{i}:\Gamma(L_{i})\to\mathfrak{X}(P)$ be the infinitesimal actions uniquely determined by the condition $\mathscr{a}_{i}\left(\mathrm{d}\mu_{i}(v),\alpha\right)=v$, where $v\in\mathfrak{X}(P)$ and $(\mathrm{d}\mu_{i}(v),\alpha)\in\Gamma(L_{i})$.
@@ 4.3 Global and infinitesimal Morita equivalences of compatible structures
The Dirac manifolds $(M_{1},L_{1},\eta_{1})$ and $(M_{2},L_{2},\eta_{2})$ are said to be Morita equivalent if there exists a manifold $P$ equipped with a $2$-form $\varpi\in\Omega^{2}(P)$, and two surjective submersions
@@
Let $(M_{1},L_{1},\eta_{1})$ and $(M_{2},L_{2},\eta_{2})$ be Morita equivalent Dirac structures. We use the same notation $L_{1}$ and $L_{2}$ for their Lie algebroids. Condition $(a)$ in Definition 4.11 implies that the map
$\mathscr{a}:\Gamma(L_{1})\oplus\Gamma(L_{2})\to\mathfrak{X}(P),\quad(\widetilde{v}_{1},\alpha_{1})\oplus(\widetilde{v}_{2},\alpha_{2})\mapsto v,$ (50)
where $v\in\mathfrak{X}(P)$ is such that $\mathrm{d}\mu_{1}(v)=\widetilde{v}_{1}$, $\mathrm{d}\mu_{2}(v)=\widetilde{v}_{2}$ and $\varpi^{\flat}(v)=\mu_{1}^{*}\alpha_{1}-\mu_{2}^{*}\alpha_{2}$, is well-defined. Define the actions
@@
by restricting the map $\mathscr{a}$ of equation (50) to $\Gamma(L_{1})\oplus\{0\}$ and $\{0\}\oplus\Gamma(L_{2})$, respectively. Since $\mu(p)=(\mu_{1}(p),\mu_{2}(p))$ is a forward Dirac map, we have that for any $\xi\oplus 0=(\mathrm{d}\mu_{1}(v_{1}),\alpha)\oplus 0\in\Gamma(L_{1})\oplus\{0\}$, $\mathscr{a}_{1}(\xi)\in\operatorname{ker}\mathrm{d}\mu_{2}$. Similarly, for any section $\zeta=(\mathrm{d}\mu_{2}(v_{2}),\alpha_{2})$ of $L_{2}$ we have that $\mathscr{a}_{2}(\zeta)\in\operatorname{ker}\mathrm{d}\mu_{1}$. From condition (b) in Definition 4.11, it follows that $\omega(v_{1},v_{2})=0$; hence, a computation of the Courant bracket $\llbracket-,-\rrbracket_{\eta}$ yields,
@@
Given that $\mu_{1}$ and $\mu_{2}$ are forward Dirac maps, we have that $\iota_{v_{1}}\varpi=\mu_{1}^{*}\alpha_{1}$ and $\iota_{v_{2}}\varpi=-\mu_{2}^{*}\alpha_{2}$. Using again $v_{1}\in\operatorname{ker}\mathrm{d}\mu_{2}$ and $v_{2}\in\operatorname{ker}\mathrm{d}\mu_{1}$, we see that $\iota_{v_{1}}\mathrm{d}\iota_{v_{2}}\varpi-\iota_{v_{2}}\mathrm{d}\iota_{v_{1}}\varpi=-\iota_{v_{1}}\mu^{*}_{2}\mathrm{d}\alpha_{2}-\iota_{v_{2}}\mu_{1}^{*}\mathrm{d}\alpha_{1}=0$. Because $L_{\varpi}$ is closed under the $\eta$-Courant bracket, it follows that
$\varpi^{\flat}\left(\left[v_{1},v_{2}\right]\right)=0,$
i.e. $\left[v_{1},v_{2}\right]\in\operatorname{ker}\varpi$. Moreover, the naturality of the commutator of vector fields implies that $\left[v_{1},v_{2}\right]$ is also in the kernel of $\mathrm{d}\mu_{1}$ and $\mathrm{d}\mu_{2}$. Using condition (a) in Definition 4.11 we obtain
$\left[\mathscr{a}_{1}(\xi),\mathscr{a}_{2}(\zeta)\right]=0,$
this is, the actions $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ commute as required by Definition 3.8.
Furthermore, the actions are injective as consequence of the forward Dirac map condition: if $\mathscr{a}_{1}(\xi)=0$ then $\rho(\xi)=\mathrm{d}\mu_{1}(0)=0$ and $(\mathrm{d}\mu_{1})^{*}\alpha=\varpi^{\flat}(0)=0$; since $\mathrm{d}\mu_{1}$ is surjective, $\alpha=0$ so $\xi=0$. The fact that $\mathscr{a}_{1}\left(\mu_{1}^{*}L_{1}\right)_{p}=\operatorname{ker}\mathrm{d}\mu_{2\,p}$ for every $p\in P$, follows because $L_{\varpi}$ determines the Dirac structure on $M_{1}\times M_{2}$ (since $\mu:P\to M_{1}\times M_{2}$ is a forward Dirac map), and hence for any $v\in\operatorname{ker}\mathrm{d}\mu_{2\,p}$ there exists $\alpha\in T_{\mu_{1}(p)}^{*}M_{1}$ such that $\mu_{1}^{*}\xi\coloneqq(\mathrm{d}\mu_{1}(v),\alpha)\in L_{1,\mu_{1}(p)}$. Therefore $\mathscr{a}_{1}\left(\mu^{*}\xi\right)=v$ as we wanted to prove, and the same argument shows that $\mathscr{a}_{2}\left(\mu_{2}^{*}L_{2}\right)_{p}=\operatorname{ker}\mathrm{d}\mu_{1\,p}$. Considering the fact that the fibers are connected and simply-connected, and the actions complete, we conclude the proof of the proposition. ∎
@@
where $\pi_{\varpi}^{\sharp}$ is the inverse of $\varpi^{\flat}:TM\to T^{*}M$. Then $(M_{1},L_{1},r_{1})$ and $(M_{2},L_{2},r_{2})$ are Morita equivalent Dirac-Nijenhuis structures if and only if there is a Morita equivalence of Poisson manifolds $(M_{1},\pi_{1})\xleftarrow{\ \mu_{1}\ }(P,\varpi)\xrightarrow{\ \mu_{2}\ }(M_{2},-\pi_{2})$ together with a Nijenhuis tensor field $J\in\Omega^{1}(P,TP)$ compatible with $\varpi$, satisfying the Morita equivalence of the corresponding 1-derivations.

### v2

@@ 4.1. Quasi-symplectic-Nijenhuis groupoids
Let $(P_{1},\varpi_{1},J_{1})$ and $(P_{2},\varpi_{2},J_{2})$ be Morita $(\mathscr{G},\mathscr{H})$- and $(\mathscr{H},\mathscr{G}^{\prime})$-bibundles, respectively, realizing Morita equivalences between the quasi-symplectic–Nijenhuis groupoids $(\mathscr{G},\omega_{1},\eta_{1},N_{1})$, $(\mathscr{H},\omega_{2},\eta_{2},N_{2})$, and $(\mathscr{G}^{\prime},\omega_{3},\eta_{3},N_{3})$. Let $P=P_{1}\diamond P_{2}$ denote the quotient of the fiber product $P_{1}\times_{H_{0}}P_{2}$ by the diagonal action of $\mathscr{H}$, and let $\pi:P_{1}\times_{H_{0}}P_{2}\to P$ be the natural projection. The proof relies on the following results:
  1. (1)
    Theorem 2.11 establishes that $(P,J_{1}\diamond J_{2})$ constitutes a Morita bibundle between the Nijenhuis groupoids $(\mathscr{G},N_{1})$ and $(\mathscr{G}^{\prime},N_{3})$, where the Nijenhuis tensor field $J_{1}\diamond J_{2}\in\Omega^{1}(P,TP)$ is the projection of $J_{1}\times J_{2}$.
  2. (2)
    By [undefam, Theorem 4.5], the pair $(P,\varpi)$ defines a Morita bibundle between the quasi-symplectic groupoids $(\mathscr{G},\omega_{1},\eta_{1})$ and $(\mathscr{G}^{\prime},\omega_{3},\eta_{3})$, where the 22-form $\varpi\in\Omega^{2}(P)$ is defined by Equation (41).
It remains only to prove the compatibility between the $(1,1)$-tensor field $J_{1}\diamond J_{2}$ and the 22-form $\varpi$ on the quotient manifold $P$.
Compatibility on the fiber product. First, we show that the compatibility holds for the $(1,1)$-tensor field $J_{1}\times J_{2}\in\Omega^{1}\left(P_{1}\times_{H_{0}}P_{2},T(P_{1}\times P_{2})\right)$ and the 22-form $\pi^{*}\varpi=\mathrm{pr}_{1}^{*}\varpi_{1}+\mathrm{pr}_{2}^{*}\varpi_{2}\in\Omega^{2}\left(P_{1}\times_{H_{0}}P_{2}\right)$. To verify the first condition, Equation (42), let $(v_{1},v_{2})$ be a vector field on $P_{1}\times_{H_{0}}P_{2}$. Using the fact that $(\varpi_{1},J_{1})$ and $(\varpi_{2},J_{2})$ are compatible pairs, we have
@@
(44) $\mathrm{d}\left(\left(\pi^{*}\varpi\right)_{J_{1}\times J_{2}}\right)=\left(\mathrm{d}\left(\pi^{*}\varpi\right)\right)_{J_{1}\times J_{2}}.$
Thus, the pair $(\pi^{*}\varpi,J_{1}\times J_{2})$ is compatible.
Descent to the quotient. Finally, these compatibility relations descend to the quotient $P$. Consider a vector field $\bar{v}\in\mathfrak{X}(P)$ as the image under $\mathrm{d}\pi$ of some $\pi$-projectable vector field $v=(v_{1},v_{2})\in\mathfrak{X}\left(P_{1}\times_{H_{0}}P_{2}\right)$. Recall that $J_{1}\times J_{2}$ preserves projectable vector fields and that it is $\pi$-related to $J_{1}\diamond J_{2}$. Using (41) and evaluating at $\bar{v}=\mathrm{d}\pi(v)$, we find:
$\varpi^{\flat}\circ(J_{1}\diamond J_{2})\left(\mathrm{d}\pi(v)\right)$ $=\varpi^{\flat}\circ\mathrm{d}\pi\left(J_{1}\times J_{2}\left(v\right)\right)$
@@
$=(J_{1}\diamond J_{2})^{*}\circ\varpi^{\flat}(\mathrm{d}\pi(v)).$
Therefore $\varpi^{\flat}\circ(J_{1}\diamond J_{2})=(J_{1}\diamond J_{2})^{*}\circ\varpi^{\flat}$. Similarly, using (44) and the naturality of the exterior derivative, we obtain:
$\left(\mathrm{d}\varpi\right)_{J_{1}\diamond J_{2}}\left(\mathrm{d}\pi(u),\mathrm{d}\pi(v),\mathrm{d}\pi(w)\right)$ $=\mathrm{d}\varpi\left(J_{1}\diamond J_{2}\left(\mathrm{d}\pi(u)\right),\mathrm{d}\pi(v),\mathrm{d}\pi(w)\right)$
@@
$=\mathrm{d}\left(\varpi_{J_{1}\diamond J_{2}}\right)\left(\mathrm{d}\pi(u),\mathrm{d}\pi(v),\mathrm{d}\pi(w)\right).$
Therefore $\mathrm{d}(\varpi_{J_{1}\diamond J_{2}})=(\mathrm{d}\varpi)_{J_{1}\diamond J_{2}}$. This completes the proof. ∎
4.2. Twisted Dirac-Nijenhuis structures
@@
(45) $\llbracket(u,\alpha),(v,\beta)\rrbracket_{\eta}\coloneqq\left(\left[u,v\right],\mathscr{L}_{u}\beta-\iota_{v}\mathrm{d}\alpha+\iota_{u\wedge v}\eta\right),\quad\eta\in\Omega^{3}(M).$
Every twisted Dirac structure $L$ is intrinsically a Lie algebroid, where the anchor is the projection to TMTM and the bracket is the restriction of the $\eta$-Courant bracket. The conditions for when an arbitrary Lie algebroid is isomorphic to a Dirac structure are detailed in [undefi].
4.2.1. Integrating Dirac structures.
The precise correspondence between twisted Dirac structures and quasi-symplectic groupoids is established in [undefi, Theorem 2.2] and [undefi, Theorem 2.4]: an integrable Lie algebroid $A_{\mathscr{G}}$ is isomorphic to an $\eta$-twisted Dirac structure $L$ on its base if and only if its source-simply connected integration $\mathscr{G}$ admits a quasi-symplectic structure. This correspondence requires the target map to be a forward Dirac map. A map $\mu:(M,L_{M})\to(N,L_{N})$ is said to be a forward Dirac map if for every $p\in M$
(46) $(L_{N})_{\mu(p)}\coloneqq\left\{(\mathrm{d}\mu(v),\alpha)\in\mathbb{T}_{\mu(p)}N\mid v\in T_{p}M,(v,(\mathrm{d}\mu)^{*}(\alpha))\in L_{M}\right\}.$
4.2.2. Twisted Dirac-Nijenhuis structures
@@
We recall from [undefh] a generalization of the compatibility for pairs given in Equations (42) and (43). Let $L\subseteq TM\oplus T^{*}M$ be a subbundle. A $(1,1)$-tensor field $r\in\Omega^{1}(M,TM)$ is said to be compatible with $L$ if
(47) $\left(r,r^{*}\right)(L)\subseteq L\qquad\text{and}\qquad\mathbb{D}^{r}_{v}\left(\Gamma(L)\right)\subseteq\Gamma(L),$
@@
Example 4.8 (Poisson-Nijenhuis manifold [undefac, undefw]).
@@
(48) $\pi^{\sharp}\circ r^{*}=r\circ\pi^{\sharp}$
(49) $R_{\pi}^{r}\left(v,\alpha\right)\coloneqq\pi^{\sharp}\left(\mathscr{L}_{v}r^{*}(\alpha)-\mathscr{L}_{r(v)}\alpha\right)-\left(\mathscr{L}_{\pi^{\sharp}(\alpha)}r\right)(v)=0,$
@@
The relation between $(1,1)$-tensor fields $r\in\Omega^{1}(M,TM)$ compatible with a Dirac structure $L\subseteq\mathbb{T}M$ and the Lie algebroid defined by $L$ is provided by [undefh, Lemma 6.1]: if $r\in\Omega^{1}(M,TM)$ is compatible with $L$, then the 11-derivation $\left(\mathbb{D}^{r}|_{\Gamma(L)},(r,r^{*}),r\right)$ is a IM 11-derivation on the Lie algebroid $L$. Moreover, if $r$ is Nijenhuis then the corresponding 11-derivation is Nijenhuis.
@@
See [undefh, Section 3.3]. ∎
4.2.3. Integrating Dirac-Nijenhuis structures.
The following theorem is an adaptation of [undefh, Theorem 6.3] to the twisted case where $\eta\neq 0$. The proof is similar.
@@
Let $(\mathscr{G},\omega,\eta)$ be a quasi-symplectic groupoid integrating the $\eta$-twisted Dirac structure $L\subseteq\mathbb{T}M$. Let $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ be a Nijenhuis $(1,1)$-tensor field integrating the 11-derivation $\left(D,\ell,r\right)$. If $\omega$ and NN are compatible (Equations (42) and (43)), then $L$ and $r$ are compatible in the sense of Equations (47), and the converse holds when $\mathscr{G}$ is source connected.
@@
4.3. Global and infinitesimal Morita equivalences of compatible structures
In this subsection, we introduce a notion of Morita equivalence for Dirac-Nijenhuis structures which corresponds to the infinitesimal counterpart of Morita equivalence for quasi-symplectic Nijenhuis groupoids. To do so, we first need to review the Morita equivalence of Dirac manifolds themselves. Similarly to how a Morita equivalence of Poisson manifolds can be expressed in terms of symplectic realizations, Morita equivalences of Dirac structures are expressed in terms of the so-called Dirac realizations, or more precisely a dual pair, a concept detailed in [undefq].
In the following, we fix two twisted Dirac manifolds $(M_{1},L_{1},\eta_{1})$ and $(M_{2},L_{2},\eta_{2})$ and consider a manifold $P$ equipped with a 22-form $\varpi\in\Omega^{2}(P)$. Consider surjective submersions $\mu_{1}:P\to M_{1}$ and $\mu_{2}:P\to M_{2}$, define $\eta\coloneqq\mu_{2}^{*}\eta_{2}-\mu_{1}^{*}\eta_{1}$, and equip $P$ with the $\eta$-twisted Dirac structure $L_{\varpi}$. We denote the opposite Dirac structure on $M_{2}$ by $\bar{L}_{2}\coloneqq\{(v,-\alpha)\mid(v,\alpha)\in L_{2}\}$. Furthermore, let $\mathscr{a}_{i}:\Gamma(L_{i})\to\mathfrak{X}(P)$ be the infinitesimal actions uniquely determined by the condition $\mathscr{a}_{i}\left(\mathrm{d}\mu_{i}(v),\alpha\right)=v$, where $v\in\mathfrak{X}(P)$ and $(\mathrm{d}\mu_{i}(v),\alpha)\in\Gamma(L_{i})$.
@@ 4.3. Global and infinitesimal Morita equivalences of compatible structures
The Dirac manifolds $(M_{1},L_{1},\eta_{1})$ and $(M_{2},L_{2},\eta_{2})$ are said to be Morita equivalent if there exists a manifold $P$ equipped with a 22-form $\varpi\in\Omega^{2}(P)$, and two surjective submersions
@@
Let $(M_{1},L_{1},\eta_{1})$ and $(M_{2},L_{2},\eta_{2})$ be Morita equivalent Dirac structures. We also denote their respective Lie algebroids by $L_{1}$ and $L_{2}$. Condition (c) in Definition 4.11 guarantees a well-defined map
(50) $\mathscr{a}:\Gamma(L_{1})\oplus\Gamma(L_{2})\to\mathfrak{X}(P),\quad(\widetilde{v}_{1},\alpha_{1})\oplus(\widetilde{v}_{2},\alpha_{2})\mapsto v,$
where $v\in\mathfrak{X}(P)$ is the unique vector field such that $\mathrm{d}\mu_{1}(v)=\widetilde{v}_{1}$, $\mathrm{d}\mu_{2}(v)=\widetilde{v}_{2}$ and $\varpi^{\flat}(v)=\mu_{1}^{*}\alpha_{1}-\mu_{2}^{*}\alpha_{2}$. We define the actions
@@
by restricting the map $\mathscr{a}$ to $\Gamma(L_{1})\oplus\{0\}$ and $\{0\}\oplus\Gamma(L_{2})$, respectively. Because $\mu=(\mu_{1},\mu_{2})$ is a forward Dirac map, any section $\xi=(\mathrm{d}\mu_{1}(v_{1}),\alpha)\in\Gamma(L_{1})$ satisfies $\mathscr{a}_{1}(\xi)\in\operatorname{ker}\mathrm{d}\mu_{2}$. Similarly, we have $\mathscr{a}_{2}(\zeta)\in\operatorname{ker}\mathrm{d}\mu_{1}$ for any section $\zeta=(\mathrm{d}\mu_{2}(v_{2}),\alpha_{2})$ of $L_{2}$.
From condition (d) in Definition 4.11, it follows that $\varpi(v_{1},v_{2})=0$. Consequently, evaluating the $\eta$-Courant bracket $\llbracket-,-\rrbracket_{\eta}$ yields
@@
Because $\mu_{1}$ and $\mu_{2}$ are forward Dirac maps, we have that $\iota_{v_{1}}\varpi=\mu_{1}^{*}\alpha_{1}$ and $\iota_{v_{2}}\varpi=-\mu_{2}^{*}\alpha_{2}$. Using the fact that $v_{1}\in\operatorname{ker}\mathrm{d}\mu_{2}$ and $v_{2}\in\operatorname{ker}\mathrm{d}\mu_{1}$, we observe that $\iota_{v_{1}}\mathrm{d}\iota_{v_{2}}\varpi-\iota_{v_{2}}\mathrm{d}\iota_{v_{1}}\varpi=-\iota_{v_{1}}\mu^{*}_{2}\mathrm{d}\alpha_{2}-\iota_{v_{2}}\mu_{1}^{*}\mathrm{d}\alpha_{1}=0$. Since $L_{\varpi}$ is closed under the $\eta$-Courant bracket, it follows that $\varpi^{\flat}\left(\left[v_{1},v_{2}\right]\right)=0$, meaning $\left[v_{1},v_{2}\right]\in\operatorname{ker}\varpi$. Furthermore, the naturality of the Lie bracket of vector fields ensures that $\left[v_{1},v_{2}\right]$ also lies in the kernels of both $\mathrm{d}\mu_{1}$ and $\mathrm{d}\mu_{2}$. Applying condition (c) in Definition 4.11, we obtain
$\left[\mathscr{a}_{1}(\xi),\mathscr{a}_{2}(\zeta)\right]=0;$
that is, the actions $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ commute, as required by Definition 3.8.
Furthermore, the actions are injective as a consequence of the forward Dirac map condition: if $\mathscr{a}_{1}(\xi)=0$ then $\rho_{1}(\xi)=\mathrm{d}\mu_{1}(0)=0$ and $(\mathrm{d}\mu_{1})^{*}\alpha=\varpi^{\flat}(0)=0$. Since $\mathrm{d}\mu_{1}$ is surjective, $\alpha=0$, and thus $\xi=0$. An analogous argument shows that $\mathscr{a}_{2}(\zeta)=0$ implies $\zeta=0$. Finally, because $\mu:P\to M_{1}\times M_{2}$ is a forward Dirac map, the identity $\mathscr{a}_{1}\left(\mu_{1}^{*}L_{1}\right)_{p}=\operatorname{ker}\mathrm{d}\mu_{2\,p}$ holds for every $p\in P$. To see this, note that for any $v\in\operatorname{ker}\mathrm{d}\mu_{2\,p}$, there exists $\alpha\in T_{\mu_{1}(p)}^{*}M_{1}$ such that $\mu_{1}^{*}\xi\coloneqq(\mathrm{d}\mu_{1}(v),\alpha)\in L_{1,\mu_{1}(p)}$. Therefore, $\mathscr{a}_{1}\left(\mu^{*}\xi\right)=v$ as desired. An identical argument establishes that $\mathscr{a}_{2}\left(\mu_{2}^{*}L_{2}\right)_{p}=\operatorname{ker}\mathrm{d}\mu_{1\,p}$. Because the fibers are connected and simply-connected, and the actions complete, the proposition follows. ∎
@@
where $\pi_{\varpi}^{\sharp}$ is the negative inverse of $\varpi^{\flat}:TM\to T^{*}M$. Then $(M_{1},L_{1},r_{1})$ and $(M_{2},L_{2},r_{2})$ are Morita equivalent Dirac-Nijenhuis structures if and only if there is a Morita equivalence of Poisson manifolds $(M_{1},\pi_{1})\xleftarrow{\ \mu_{1}\ }(P,\varpi)\xrightarrow{\ \mu_{2}\ }(M_{2},-\pi_{2})$ together with a Nijenhuis tensor field $J\in\Omega^{1}(P,TP)$ compatible with $\varpi$, satisfying the Morita equivalence of the corresponding 1-derivations.